OpenStudy (anonymous):

a round lake has a circular shoreline that is 16 miles in diameter find the are of the lake give your answer to the nearest hundredth

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

what do you know about the area of a circle of radius R?

OpenStudy (anonymous):

radius is half of the diameter

OpenStudy (anonymous):

right. so what is the radius in your problem?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

right. so the radius is 8 miles. what is the formula for area of a circle?

OpenStudy (anonymous):

A=PIE R SQUARE

OpenStudy (anonymous):

okay good. so what is the area of the circular lake with radius 8 miles?

OpenStudy (anonymous):

201.06192

OpenStudy (anonymous):

22/7 x 8 x 8 = approx 200 mil squared

OpenStudy (anonymous):

\[A = \pi (8)^{2} = \pi \times 64 = 201.06193 sq.miles\]

OpenStudy (anonymous):

okay very good. now round it off to the nearest hundredth.

OpenStudy (anonymous):

200

OpenStudy (anonymous):

no no. what is one tenth?

OpenStudy (anonymous):

NO IDEA

OpenStudy (anonymous):

one-tenth is 0.1 similarly one-hundredth is 0.01. so when some one says round it off to the nearest hundredth, they mean round it off to the nearest one hundredth. so the nearest hundredth is 201.06 sq. miles